# Optimal Convex Cover as Collision-free Space Approximation for Trajectory Generation

### v1

@@ Abstract
We propose an online iterative algorithm to find a suitable convex cover to under-approximate the free space for autonomous navigation to delineate Safe Flight Corridors (SFC). The convex cover consists of a set of polytopes such that the union of the polytopes represents obstacle-free space, allowing us to find trajectories for robots that lie within the convex cover. In order to find the SFC that facilitates optimal trajectory generation, we iteratively find overlapping polytopes of maximum volumes that include specified waypoints initialized by a geometric or kinematic planner. Constraints at waypoints appear in two alternating stages of a joint optimization problem, which is solved by a method inspired by the Alternating Direction Method of Multipliers (ADMM) with partially distributed variables. We validate the effectiveness of our proposed algorithm using a range of parameterized environments and show its applications for two-stage motion planning.
Keywords:
Motion and Path Planning, Computational Geometry, Collision Avoidance
1 Introduction
@@ 2 Related works
3 Problem Statement
3.1 Trajectory Optimization Formulation
@@
3.2 Convex Cover Generation
@@
$\Gamma(p)=\{x\in\mathbb{R}^{n}|\ x=\theta p_{i-1}+(1-\theta)p_{i},\ \forall\theta\in[0,1],\ \forall i={1,\cdots,M}\}.$ (2)
@@
3.3 Trajectory Optimization with SFC
@@
$\sigma(t)=\{\sigma_{i}(t)\ |\ t\in[\sum_{k=1}^{i-1}\Delta t_{k},\sum_{k=1}^{i}\Delta t_{k}],\ \tau=[\Delta t_{1},\cdots,\Delta t_{M}]^{T}\}.$ (5)
@@ 4 Proposed Algorithms
$+\frac{\rho}{2}\sum_{i=1}^{M}\|[g(h_{0}(\mathcal{E}_{i},p_{i})),g(h_{0}(\mathcal{E}_{i},p_{i-1}))]^{T}+y_{i}\|_{2}^{2}-\frac{\rho}{2}\|y\|_{2}^{2},$
@@
4.1 Initialization and Polytopes Generation
Given an initial path $p^{0}$, we upsample waypoints by adding auxiliary intermediate waypoints between consecutive original waypoints to avoid creating excessively long ellipsoids that exceed a specified distance threshold $\alpha$. We use the procedure InitEllipsoids that takes as input line segments of the initial path and outputs $L_{i}(L_{i})^{T}=R_{i}D_{i}(R_{i})^{T}$, where $D_{i}={\rm diag}\{\|p_{i}^{0}-p_{i-1}^{0}\|_{2}^{2}/4,\epsilon^{2},\epsilon^{2}\},R_{i}\in\mathbb{R}^{n}$ is the rotation from the orientation of line vector. $\epsilon$ is a constant, usually chosen to be smaller than the map resolution. The geometric path is assumed to be collision-free, which guarantees the collision-free of initial ellipsoids.
In InflatePolytopes, to find a maximum volume of polytopes in $\mathcal{X}_{free}$, we reformulate the optimization problem (4) as
@@ 4.1 Initialization and Polytopes Generation
We use a similar method in [3] to solve this problem by finding tangent separating planes from the ellipsoid. To save computation, we define a local bounding box by adding a margin of ±$\pm l$ around the ellipsoid for each polytope generation. This margin ensures that all local obstacles within this expanded range are considered. Hence, the overlapping condition $\mathcal{P}_{i}\cap\mathcal{P}_{i+1}\neq\varnothing$ is naturally satisfied by waypoint-constrained ellipsoids. Additionally, we store the obstacle points near each initialized ellipsoid for parallel updating of polytopes during iteration.
4.2 Waypoints Update with Trajectory Heuristics
@@
$\vspace{-0.3cm}\min_{p,\tau}$ $w_{c}\tilde{J}(p,\tau)+\sum_{i=1}^{M}\frac{\rho}{2}\|[g(h_{0}(\mathcal{E}_{i}^{k},p_{i})),g(h_{0}(\mathcal{E}_{i}^{k},p_{i-1}))]^{T}+y_{i}\|_{2}^{2},$ (11a)
@@
4.2.1 Minimum Path Length.
@@
$\min_{p}$ $w_{c}\sum_{i=1}^{M}\|p_{i}-p_{i-1}\|_{2}+\sum_{i=1}^{M}\frac{\rho}{2}\|[g(h_{0}(\mathcal{E}_{i}^{k},p_{i})),g(h_{0}(\mathcal{E}_{i}^{k},p_{i-1}))]^{T}+y_{i}^{k}\|_{2}^{2},$ (12a)
${\rm s.t.}$ $p_{i}\in\mathcal{P}_{i}\cap\mathcal{P}_{i+1},\ \forall i=1,\cdots,M-1,$ (12b)
where constraint (12b) ensures that waypoints are generated inside the overlapping region of every two consecutive polytopes.
@@
4.2.2 Minimum Control Effort.
@@
${\rm s.t.}$ $M\sigma(p,\tau)=e,$ (13b)
$A_{i}\sigma_{i}(p,\tau)\leq b_{i},\ \forall i=1,\cdots,M,$ (13c)
where $\sigma_{i}$ is the $i$-th trajectory segment mapping from derivative vectors and time intervals, which has been discussed in [31]. The constraints (13b) include boundary and continuity constraints for polynomials, and constraints (13c) enforces trajectory inside corridors. Practically, we can use a fixed time allocation with a time factor to provide sub-optimal heuristics.
4.3 Maximum Volume Ellipsoids with Inner&Outer Constraints
@@
$\min_{L_{i},d_{i}}$ $-w_{v}\log\det(L_{i})+\frac{\rho}{2}\|[g(h_{0}(\mathcal{E}_{i},p_{i}^{k})),g(h_{0}(\mathcal{E}_{i},p_{i-1}^{k}))]^{T}+y_{i}^{k}\|_{2}^{2},$ (14a)
${\rm s.t.}$ $\|(a_{i,j})^{T}L_{i}\|_{2}+(a_{i,j})^{T}d_{i}\leq b_{i,j},\ \ j=1,\cdots,v_{i}$ (14b)
$L_{i}\ {\rm is\ lower\ triangular},$ (14c)
where the first term of objective, $\log\det(L_{i})$, is proportional to the volume of the ellipsoid, and (14b) represents corridor constraints. We employ a distributed style to update $\mathcal{E}^{k+1}$.
5 Results
5.1 Implementation Details
We use the procedure developed in [26] which defines a quantity called the Environmental Complexity Signature (ECS) to generate environments of varying complexities (obstacle structures, densities, etc). All environments are stored as voxel maps. We use the RRT* [2] algorithm from the OMPL library [32] to get the initial geometric path for all methods. We use the L-BFGS algorithm [33] to minimize the augmented Lagrangian for all the constrained optimization problems in our approach. We therefore replace hard constraints with soft ones, but we do not perform a dual update step in our approach. In practice, we run the inner iteration once and update the polytope through several outer iterations. Our approach can be deployed on our custom hardware platform [34] to provide an environmental representation for two-stage motion planning. [^1] We will open-source code at https://github.com/KumarRobotics/kr˙opt˙sfc.
5.2 Numerical Analysis
@@
5.2.1 Ellipsoids and Polytopes over Iterations.
@@
5.2.2 The Effects of Cost Heuristics.
@@
Figure 8: Performance by different distinct settings with $l\in\{1.5,2\}$ and $\alpha\in\{2,3\}$.
5.2.3 Configurable Performance in Different Environments
@@
We leverage parameters such as local range $l$ and upsample threshold $\alpha$ to generate SFC with specified accuracy and efficiency. Our analysis considers the following four setups with a local range of 1.5 m and 2.0 m and an upsample threshold of 2.0 m and 3.0 m. The upsample threshold influences the number of segments, and the local range determines the number of obstacles considered for each polytope generation. These settings are evaluated in terms of computation time, trajectory cost ($J$), number of segments ($M$), path length ($\tilde{J}_{1}$), and volumes of ellipsoids, polytopes, and overlapping areas. Trajectory cost is evaluated by optimizing time intervals and waypoints in [35]. As shown in Fig.8, the computation time can be reduced by decreasing the local range and increasing the upsample threshold. The path length, polytope volumes, overlapping volumes, and trajectory cost also decrease accordingly. We conclude that within the same homotopic class, the optimal trajectory is indeed influenced by the convex cover.
5.3 Comparison of Convex Cover
@@
We compare our proposed algorithm with two state-of-the-art SFC generation methods, IRIS[3] and the method presented in [8], denoted as RILS. IRIS was originally designed for graph cover, and we use its applications for a convex cover of a path with incremental point collision checking. The iteration number of IRIS is set as the major iteration in their paper, the same as our proposed method. To get inner Löwner-John ellipsoids in IRIS, we directly employ the Cholesky representation for optimization in [7]. We evaluate the performance of three methods across various randomized environments with dimensions of $30\times 30\times 5$m3 characterized by the ECS (density index, cluster index, and structure index). We conduct 10000 tests within these environments, randomly selecting start and end points with a distance threshold of 5 m. The results are filtered to exclude all invalid cases encountered by any of the methods.
Figure 9: Benchmark comparisons with increasing density index (defined in [26]).
@@
We categorize test cases based on the range of the density index to represent environment complexity and show the results in Fig. 9. As the density index increases, environments become more complex, which leads to reduced polytope volumes and decreased trajectory performance across all three methods. When the density index exceeds 0.5, finding a feasible geometric path over long distances becomes challenging, resulting in shorter paths and fewer segments. Our method updates the path for better trajectory generation, particularly in complex environments such as narrow mazes where RILS show poor performance around corners, illustrated in Fig. 10.
Figure 10: The illustration of SFC generated by three methods.
@@
6 Conclusion and Future Work
This paper introduces a novel algorithm to find an optimal convex cover of collision-free space as SFC for trajectory optimization.
Figure 11: Benchmark comparisons with increasing start-end distance.
The convex cover is formulated as a joint optimization of ellipsoids, polytopes, and intermediate waypoints, which are updated iteratively to find optimal solutions.
The proposed method can be applied to two-stage motion planning as an intermediate process, or optimizing with trajectory generation jointly. Furthermore, we will extend our method to whole-body SE(n) trajectory planning with an initial set of rigid bodies.
Acknowledgements
We gratefully acknowledge the support of The Institute for Learning-Enabled Optimization at Scale (TILOS) funded by the National Science Foundation (NSF) under NSF grants CCR-2112665, USDA/NIFA 2022-67021-36856, IoT4Ag ERC funded through NSF Grant EEC-1941529, and ONR grant N00014-20-S-B001.

### v2

@@ Abstract
We propose an online iterative algorithm to find an optimal convex cover to under-approximate the free space for autonomous navigation to delineate Safe Flight Corridors (SFC). The convex cover consists of a set of polytopes such that the union of the polytopes represents obstacle-free space, allowing us to find trajectories for robots that lie within the convex cover. In order to find the SFC that facilitates optimal trajectory generation, we iteratively find overlapping polytopes of maximum volumes that include specified waypoints initialized by a geometric or kinematic planner. Constraints at waypoints appear in two alternating stages of a joint optimization problem, which is solved by a method inspired by the Alternating Direction Method of Multipliers (ADMM) with partially distributed variables. We validate the effectiveness of our proposed algorithm using a range of parameterized environments and show its applications for two-stage motion planning.
I Introduction
@@ II Related works
III Problem Statement
III-A Trajectory Optimization Formulation
@@
III-B Convex Cover Generation
@@
$\Gamma(p)=\{x\in\mathbb{R}^{n}|\ x=\theta p_{i-1}+(1-\theta)p_{i},$ (2)
$\forall\theta\in[0,1],\ \forall i={1,\cdots,M}\}.$
@@
III-C Trajectory Optimization with SFC
@@
$\sigma(t)=\{\sigma_{i}(t)\ |\ t\in[\sum_{k=1}^{i-1}\Delta t_{k},\sum_{k=1}^{i}\Delta t_{k}],$ (5)
$\quad\ \tau=[\Delta t_{1},\cdots,\Delta t_{M}]^{T}\}.$
@@ IV Proposed Algorithms
$+\frac{\rho}{2}\sum_{i=1}^{M}\|[g(h_{0}(\mathcal{E}_{i},p_{i})),g(h_{0}(\mathcal{E}_{i},p_{i-1}))]^{T}$
$+y_{i}\|_{2}^{2}-\frac{\rho}{2}\|y\|_{2}^{2},$
@@
IV-A Initialization and Polytopes Generation
Given an initial path $p^{0}$, we upsample waypoints by adding auxiliary intermediate waypoints between consecutive original waypoints to avoid creating excessively long ellipsoids that exceed a specified distance threshold $\alpha$. We use the procedure InitEllipsoids that takes as input line segments of the initial path and outputs the ellipsoid with $d_{i}=(p_{i}+p_{i-1})/2$ and
$L_{i}L_{i}^{T}=v_{i}v_{i}^{T}\frac{||p_{i}-p_{i-1}||_{2}^{2}}{4}+(I-v_{i}v_{i}^{T})\epsilon^{2}.$ (10)
We assume the initial path has $\epsilon$-clearance from obstacles; this implies that the constructed ellipsoids are collision-free. In InflatePolytopes, to find a maximum volume of polytopes in $\mathcal{X}_{free}$, we reformulate the optimization problem (4) as
@@ IV-A Initialization and Polytopes Generation
We use a similar method in [3] to solve this problem by finding tangent separating planes from the ellipsoid. To save computation, we define a local bounding box by adding a margin of ±$\pm l$ around the ellipsoid for each polytope generation. This margin ensures that only the local obstacles within this expanded range are considered. The overlap condition $\mathcal{P}_{i}\cap\mathcal{P}_{i+1}\neq\varnothing$ is satisfied by construction. Additionally, we store the obstacle points near each initialized ellipsoid for parallel updating of polytopes during iteration.
IV-B Waypoints Update with Trajectory Heuristics
@@
$\vspace{-0.3cm}\min_{p,\tau}$ $w_{c}\tilde{J}(p,\tau)+$ (12a)
$\sum_{i=1}^{M}\frac{\rho}{2}\|[g(h_{0}(\mathcal{E}_{i}^{k},p_{i})),g(h_{0}(\mathcal{E}_{i}^{k},p_{i-1}))]^{T}+y_{i}\|_{2}^{2},$
@@
IV-B1 Minimum Path Length.
@@
$\min_{p}$ $w_{c}\sum_{i=1}^{M}\|p_{i}-p_{i-1}\|_{2}+$ (13a)
$\sum_{i=1}^{M}\frac{\rho}{2}\|[g(h_{0}(\mathcal{E}_{i}^{k},p_{i})),g(h_{0}(\mathcal{E}_{i}^{k},p_{i-1}))]^{T}+y_{i}^{k}\|_{2}^{2},$ (13b)
${\rm s.t.}$ $p_{i}\in\mathcal{P}_{i}\cap\mathcal{P}_{i+1},\ \forall i=1,\cdots,M-1,$ (13c)
where constraint (13c) ensures that waypoints are generated inside the overlapping region of every two consecutive polytopes.
@@
IV-B2 Minimum Control Effort.
@@
$\sum_{i=1}^{M}\frac{\rho}{2}\|[g(h_{0}(\mathcal{E}_{i}^{k},p_{i})),g(h_{0}(\mathcal{E}_{i}^{k},p_{i-1}))]^{T}+y_{i}^{k}\|_{2}^{2},$ (14b)
${\rm s.t.}$ $M\sigma(p,\tau)=e,$ (14c)
$A_{i}\sigma_{i}(p,\tau)\leq b_{i},\ \forall i=1,\cdots,M,$ (14d)
where $\sigma_{i}$ is the $i$-th trajectory segment mapping from derivative vectors and time intervals, which has been discussed in [31]. The constraints (14c) include boundary and continuity constraints for polynomials, and constraints (14d) enforces trajectory inside corridors. Practically, we can use a fixed time allocation with a time factor to provide sub-optimal heuristics.
IV-C Maximum Volume Ellipsoids with Inner&Outer Constraints
@@
$\min_{L_{i},d_{i}}$ $-w_{v}\log\det(L_{i})+$ (15a)
$\frac{\rho}{2}\|[g(h_{0}(\mathcal{E}_{i},p_{i}^{k})),g(h_{0}(\mathcal{E}_{i},p_{i-1}^{k}))]^{T}+y_{i}^{k}\|_{2}^{2},$ (15b)
${\rm s.t.}$ $\|(a_{i,j})^{T}L_{i}\|_{2}+(a_{i,j})^{T}d_{i}\leq b_{i,j},\ \ j=1,\cdots,v_{i}$ (15c)
$L_{i}\ {\rm is\ lower\ triangular},$ (15d)
where the first term of objective, $\log\det(L_{i})$, is proportional to the volume of the ellipsoid, and (15c) represents corridor constraints. We employ a distributed style to update $\mathcal{E}^{k+1}$.
V Results
V-A Implementation Details
We employ the environments provided in  [26] which include both simulated and real-world point clouds for cluttered obstacle maps, maze maps, and real-world maps. We follow their procedure which defines a quantity called the Environmental Complexity Signature (ECS) to generate simulated environments of varying complexities (obstacle structures, densities, etc). All environments are stored as voxel maps. We use the RRT* [2] algorithm from the OMPL library [32] to get the initial geometric path for all methods. We use the L-BFGS algorithm [33] to minimize the augmented Lagrangian for all the constrained optimization problems in our approach and replace the rest hard constraints with penalties. In practice, we run the inner iteration once and update the polytope through several outer iterations. Our approach can be deployed on our custom hardware platform [34] to provide an environmental representation for two-stage motion planning.  [^1] We will open-source at https://github.com/KumarRobotics/kr˙opt˙sfc.
V-B Numerical Analysis
@@
V-B1 Ellipsoids and Polytopes over Iterations.
@@
V-B2 The Effects of Cost Heuristics.
@@
Figure 8: Performance by different settings with $l\in\{1.5,2\}$ and $\alpha\in\{2,3\}$.
V-B3 Configurable Performance in Different Environments
@@
We leverage parameters such as local range $l$ and upsample threshold $\alpha$ to generate SFC with specified accuracy and efficiency. Our analysis considers the following four setups with a local range of 1.5 m and 2.0 m and an upsample threshold of 2.0 m and 3.0 m. The upsample threshold influences the number of segments, and the local range determines the number of obstacles considered for each polytope generation. These settings are evaluated in terms of computation time, trajectory cost ($J$), number of segments ($M$), path length ($\tilde{J}_{1}$), volumes of polytopes, and overlapping areas. Trajectory cost is evaluated by optimizing time intervals and waypoints in [35]. As shown in Fig.8, the computation time can be reduced by decreasing the local range and increasing the upsample threshold. The path length, polytope volumes, overlapping volumes, and trajectory cost also decrease accordingly. We conclude that within the same homotopic class, the optimal trajectory is indeed influenced by the convex cover.
V-C Comparison of Convex Cover
@@
We compare our proposed algorithm with two state-of-the-art SFC generation methods, IRIS[3] and the method presented in [8], denoted as RILS. IRIS was originally designed for graph cover, and we use its applications for a convex cover of a path with incremental point collision checking. The iteration number of IRIS is set as the major iteration in their paper, the same as our proposed method. To get inner Löwner-John ellipsoids in IRIS, we directly employ the Cholesky representation for optimization in [7]. The local range for considering obstacles within the bounding box is set to $l=2$ m to optimize computational efficiency. We first evaluate the performance of three methods across three types of randomly sampled maps generated from [26]. We conduct 1,000 trials for each type of environment, with start and end points randomly selected, ensuring a minimum distance of 10 meters between them. As shown in Table I, our method outperforms other benchmark methods for generating better trajectories, specifically in unstructured environments.
TABLE I: Comparison on Different Types of Maps.
Map Method
$J$
($m^{2}/s^{5}$)
$vol(\mathcal{P})$
($m^{3}$)
Time ($ms$) $\tilde{J}_{1}$ ($m$)
Obstacle IRIS* 35.39 107.82 45.93 20.17
RILS 45.76 60.08 11.52 19.37
Ours 23.83 84.79 24.76 19.66
Maze IRIS* 33.82 118.38 33.85 20.31
RILS 33.48 83.93 6.06 18.81
Ours 33.02 121.79 30.46 19.56
Real IRIS* 20.53 144.86 17.68 14.04
RILS 28.28 158.45 3.11 13.80
Ours 16.45 178.72 19.32 14.04
*IRIS search range is cropped.
We visualize one trial as an instance, illustrated in Fig. 9. Our method iteratively refines the path for better trajectory generation, particularly in complex environments such as narrow mazes where RILS show poor performance around corners.
Figure 9: The illustration of SFC generated by three methods. More comparisons are available at https://www.youtube.com/watch?v=J3YmIk_TKiQ
To further evaluate the performance, we conduct 10,000 additional tests in a broader set of simulated environments with dimensions of $30\times 30\times 5$m3 characterized by the ECS (density index, cluster index, and structure index). The results are filtered to exclude all invalid cases encountered by any of the methods. We categorize test cases based on the range of the density index to represent environment complexity and show the results in Fig. 10.
Figure 10: Comparisons with increasing density index (defined in [26]).
@@
As the density index increases, environments become more complex, which leads to reduced polytope volumes and decreased trajectory performance across all three methods. When the density index exceeds 0.5, finding a feasible geometric path over long distances becomes challenging, resulting in shorter paths and fewer segments.
Figure 11: Comparisons with increasing start-end distance.
@@
VI Conclusion and Future Work
This paper introduces a novel algorithm to find an optimal convex cover of collision-free space as SFC for trajectory optimization. The convex cover is formulated as a joint optimization of ellipsoids, polytopes, and intermediate waypoints, which are updated iteratively to find optimal solutions. The proposed method can be applied to two-stage motion planning as an intermediate process, or optimizing with trajectory generation jointly. Our future work will address whole-body trajectory planning for $SE$(2) and $SE$(3).